1). Bill's Roast Beef sells 5 times as many sandwiches as Pete's Deli. The difference between their sales is 360 sandwiches. How many sandwiches did each sell?
Answer by JulietG(1812)

I'm hungry!

B = 5P
B - P = 360
Substitute the known value of P from the first equation into the second.
B - P = 360
(5P) - P = 360
4P = 360
Divide each side by 4
P = 90
If Pete sells 90, then Bill sells 450 (five times as many). 450-90=360